If events A and B are independent, then P(B|A) is equal to ________.

A. P(A)

B. P(B)

C. P(A∩B)

D. P(AUB)
Answer #1

ANSWER :- Given that A and B are independent hean P(ANB) =P(A) P(B) We know that P(BNA) P(BYR) - P(A) PAT.PB) PLA = P (B Henc